write vector v=(4,9,19) as a linear combination of vectors u1=(1,-2,3),u2=(3,-7,10),u3=(2,1,9)


1
Expert's answer
2021-09-22T22:55:18-0400


Given that:

"\vec v=4\hat i +9\hat j+19\hat k, \vec u_1=\hat i-2\hat j+3\hat k,\\
\vec u_2=3\hat i-7\hat j+10\hat k,\vec u_3=2\hat i+\hat j+9\hat k\\"

Let "\vec v=x(\vec u_1)+y(\vec u_2)+z(\vec u_3)\\"

"\Rightarrow 4\hat i +9\hat j+19\hat k=x(\hat i-2\hat j+3\hat k)+y(3\hat i-7\hat j+10\hat k)+z(2\hat i+\hat j+9\hat k)\\
\Rightarrow4\hat i +9\hat j+19\hat k=(x+3y+2z)\hat i+(-2x-7y+z)\hat j+(3x+10y+9z)\hat k\\"

Comparing both sides, we get:

"x+3y+2z=4\ldots(1)\\
-2x-7y+z=9\ldots(2)\\
3x+10y+9z=19\ldots(3)\\"

Now, solving these "3" equations:

"D=\left|\begin{array}{ccc}
1 & 3 & 2 \\
-2 & -7 & 1 \\
3 & 10 & 9
\end{array}\right|=-8\\
D_{x}=\left|\begin{array}{ccc}
4 & 3 & 2 \\
9 & -7 & 1 \\
19 & 10 & 9
\end{array}\right|=-32\\
D_{y}=\left|\begin{array}{ccc}
1 & 4 & 2 \\
-2 & 9 & 1 \\
3 & 19 & 9
\end{array}\right|=16\\
D_{z}=\left|\begin{array}{ccc}
1 & 3 & 4 \\
-2 & -7 & 9 \\
3 & 10 & 19
\end{array}\right|=-24\\"

Using Cramer's rule, we get:

"\begin{gathered}
x=\frac{D_{x}}{D}, y=\frac{D_{y}}{D}, z=\frac{D_{z}}{D} \\
x=\frac{D_{x}}{D}=\frac{-32}{-8}=4 \\
y=\frac{D_{y}}{D}=\frac{16}{-8}=-2 \\
z=\frac{D_{z}}{D}=\frac{-24}{-8}=3
\end{gathered}"

"\therefore \vec v=4\vec u_1-2\vec u_2+3\vec u_3"